To compare the results of boys and girls in a class, a special test was given to 50 boys who averaged 67.4





with Sd. of 5.0, and 50 girls averaged 62.8 with Sd. of 4.6. (Group Assignment) (2 pts.)





a) Test, at





  0.05





, whether the difference is significant or not.





b) Test if the difference is 3.0.

1
Expert's answer
2022-01-28T15:11:47-0500

We have that,

Boys

"n_1=50\\\bar x=67.4\\s_1=5"

Girls

"n_2=50\\\bar x_2=62.8\\s_2=4.6"

Before we go to test the means first, we have to test their variability using F-test. This is to know whether the population variances are either equal or unequal in order to apply the appropriate test for the difference in means.

Therefore, we first test,

"H_0:\sigma_1^2=\sigma^2_2\\vs\\H_1:\sigma_1^2\not=\sigma^2_2"

The test statistic is,

"F_{cal}={s_1^2\over s_2^2}={5\over4.6}=1.086957"

The critical value is given as,

"F_{\alpha,n_1-1,n_2-1}=F_{0.05,49,49}=1.607289" and reject the null hypothesis if "F_{cal}\gt F_{0.05,49,49}"

Since

"F_{cal}=1.086957\lt F_{0.05,49,49}=1.607289", we accept the null hypothesis that the population variances for boys and girls are equal.

"a)"

We test the following hypotheses,

"H_0:\mu_1-\mu_2=0\\vs\\H_1:\mu_1-\mu_2\not=0"

We apply t distribution to perform this test as follows,.

The test statistic is given as,

"t_c={(\bar x_1-\bar x_2)-(\mu_1-\mu_2)\over \sqrt{sp^2({1\over n_1}+{1\over n_2})}}"

where "sp^2" is the pooled sample variance given as,

"sp^2={(n_1-1)s_1^2+(n_2-1)s_2^2\over n_1+n_2-2}={(49\times25)+(49\times21.16)\over98}={2261.84\over98}=23.08"

Therefore,

"t_c={(67.4-62.8)-0\over \sqrt{23.08({1\over 50}+{1\over 50})}}={4.6\over0.96083297}=4.7875"

"t_c" is compared with the table value at "\alpha=0.05" with "n_1+n_2-2=50+50-2=98" degrees of freedom.

The table value is,

"t_{{0.05\over2},98}=t_{0.025,98}=1.984467"

The null hypothesis is rejected if "|t_c|\gt t_{0.025,98}."

Now,

"|t_c|=4.7875\gt t_{0.025,98}=1.984467," and we reject the null hypothesis and conclude that there is sufficient evidence to show that the difference in means between boys and girls is significant at 5% level of significance.


"b)"

The hypothesis tested are,

"H_0:\mu_1-\mu_2=3\\vs\\H_1:\mu_1-\mu_2\not=3"

"t_c={(\bar x_1-\bar x_2)-(\mu_1-\mu_2)\over \sqrt{sp^2({1\over n_1}+{1\over n_2})}}={(67.4-62.8)-3\over\sqrt{23.08({1\over50}+{1\over 50})}}={(4.6-3)\over\sqrt{0.9232}}={1.6\over 0.96083297}=1.6652"

"t_c" is compared with the table value at "\alpha=0.05" with "n_1+n_2-2=50+50-2=98" degrees of freedom.

The table value is,

"t_{{0.05\over2},98}=t_{0.025,98}=1.984467"

The null hypothesis is rejected if "|t_c|\gt t_{0.025,98}."

Since "|t_c|=1.6652\lt t_{0.025,98}=1.984467," we fail to reject the null hypothesis and conclude that there is sufficient evidence to show that the difference in means between boys and girls is 3.